st

Showing posts with label cbse 12th mathematics. Show all posts
Showing posts with label cbse 12th mathematics. Show all posts

Monday, January 16, 2017

integral ∫ { cosx/ [ sinx +cosx] } dx

integral  ∫  { cosx / [ sinx +cosx]  } dx

first express the numerator in terms of the denominator and its derivative
split into two integrals ;one  of them is of the form u' / u which will integrate to ln(u)


 PAGE 1 BASIC INTEGRATION

PAGE 2 INTEGRATION BY SUBSTITUTION 

PAGE 3 INTEGRATION BY COMPLETION OF SQUARES

PAGE 4 INTEGRATION BY PARTS

PAGE 5 INTEGRATION BY MANIPULATION OF NUMERATOR IN TERMS OF DENOMINATOR

 disclaimer:
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work

integral of linear expression in the numerator mixed with a quadratic expression

integral of  ∫ { (4x - 7) / (x² + x +1)} dx

here there is a linear expression in the numerator mixed with quadratic expression containing both (x^2)  and x terms

try to express the linear expression in terms of the derivative of the quadratic expression and a constant.
then break the numerator into two parts.
 one will be of the form [f '(x) / f(x)] , for the second part use completion of squares.








disclaimer:
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work 

integral of 1 / sqrt[4 + 3x - x² ] using completion of squares method

integral of 1 / sqrt[4 + 3x - x² ] using completion of squares method

integrate after completing the square
to make the coefficient of x²  (+1) , take -1 common , but it remains inside the sqrt sign
to get sqrt [ - ( x² -3x - 4)] now take half the coefficient of x , that is half of (-3)
to get [x- (3/2) ]² and then adjust the constant term and finally re introduce the -1 taken  out earlier




PAGE 1 BASIC INTEGRATION

PAGE 2 INTEGRATION BY SUBSTITUTION 

PAGE 3 INTEGRATION BY COMPLETION OF SQUARES

PAGE 4 INTEGRATION BY PARTS

PAGE 5 INTEGRATION BY MANIPULATION OF NUMERATOR IN TERMS OF DENOMINATOR

disclaimer:
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work

integral using completion of squares method

integral using completion of squares method

integral of  1 / [(x^2) - x +1 ]

completion of squares method is used if there is a (x^2) term and a x term
  remember to make the coefficient of x² unity
take half the coefficient of x  that is half of  (-1)
so that you get [ x -(1/2) ]² , then adjust the constant term




PAGE 1 BASIC INTEGRATION
PAGE 2 INTEGRATION BY SUBSTITUTIONPAGE 3 INTEGRATION BY COMPLETION OF SQUARES
PAGE 4 INTEGRATION BY PARTS
PAGE 5 INTEGRATION BY MANIPULATION OF NUMERATOR IN TERMS OF DENOMINATOR




disclaimer:
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work

Friday, December 30, 2016

some more problems from binomial distribution for cbse ncert class xii mathematics probability

some more problems from binomial distribution for cbse ncert class xii  mathematics probability


4.Five cards are drawn successively with replacement from a well-shuffled deck of 52 cards. What is the probability that (i) all the five cards are spades (ii) only 3 cards are spades and (iii) none is a spade?

Let X be the number of spades among the five cards.
Assume X follows Binomial distribution with
n = 5
p = (13/52) = ( 1 / 4 ) [13 spades among the 52 cards ]
q =1 -p
q = (3 / 4)
P[X=r] = nCr prq(n-r) , r = 0,1,2,...,n
P[X=r] = 5Cr (1/4)r(3/4)(5-r) , r = 0,1,2,...,5

P[all the five cards are spades ] = P[X=5] =5C5(1/4)5(3/4)(5-5) =( 1 / 1024 )

P[only three cards are spades ] = P[X=3] =5C3(1/4)3(3/4)(5-3) =(90/1024)=(45/512)

P[none is a spade ] = P[X=0] =5C0(1/4)0(3/4)(5-0) =(243/1024)


9.On a multiple choice examination with three possible answers for each of the five questions, what is the probability that a candidate would get four or more correct answers just by guessing ?

Let X be the number of questions he answered correctly out of the 5 questions. just by guessing
Assume X follows Binomial distribution with
n =5
p = ( 1 / 3 ) [one out the three possible answers is correct and the candidate is guessing ]
q = 1 – p = ( 2 / 3 )
P[X=r] = nCr prq(n-r) , r = 0,1,2,...,n
P[X=r] = 5Cr ( 1 / 3 )r( 2 / 3 )(5 - r) r = 0, 1 ,..., 5
P[ candidate would get four or more correct answers just by guessing ] = P[X=4] + P[X=5]
= 5C4 ( 1 / 3 )4( 2 / 3 )(5 - 4) + 5C5 ( 1 / 3 )5( 2 / 3 )(5 - 5) = ( 11/243 )
=============================================================
link to index of other miscellaneous problems on probability of cbse ncert 12th mathematics

index of more problems on baye's theorem for ncert cbse mathematics 


disclaimer:
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work
 

binomial distribution problem for ncert cbse 12th mathematics probability

binomial distribution problem for ncert cbse 12th mathematics probability

1.A die is thrown 6 times. If getting an odd number is a success, what is the probability of (i) 5 successes (ii) at least 5 successes (iii) at most 5 successes?
Let X be the number of successes out of 6 throws
Assume X follows binomial distribution with
n = 6,
p = (3/6) = ( 1 / 2 )
q = 1 -p
q = ( 1 / 2 )

P[X=r] = nCr prq(n-r) , r = 0,1,2,...,n
P[X=r] = 6Cr ( 1 / 2 )r( 1 / 2 )(6 - r)

P[X=r] = 6Cr ( 1 / 2 )6

P[ 5 successes ] =P[X=5] = 6C5 ( 1 / 2 )6 = (3 /32)

P[ at least 5 successes ] =P[X=>5]= P[X=5] +P[X=6]= (7/64)

P[ at most 5 successes ] =P[X<=5] =1 - P[X=6]= 1 - 6C6 ( 1 / 2 )6 = (63 /64)
===============================
>
2.A pair of dice is thrown 4 times. If getting a doublet is considered a success, find the probability of
two successes.

Let X be the number of doublets out of 4 tosses of a pair of dice.

Assume X follows binomial distribution with
n = 4,
p = (6 / 36) [6 doublets out of 36 possible outcomes in one toss of a pair of dice]
p = ( 1 / 6)
q= 1 -p
q = (5 /6)

P[X=r] = nCr prq(n-r) , r = 0,1,2,...,n

P[X=r] = 4Cr (1/6)r(5/6)(4 - r) , r = 0,1,2,3,4
P[two successes] = P[X = 2] =4C2 (1/6)2(5/6)(4 – 2) = (25/216)
=============================================
 
 



This problem and the answer is from binomial distribution in the chapter on probability for class xii of cbse ncert 12th mathematics and is useful for the students preparing for the board examination 

disclaimer:
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work

finding n of binomial distribution given the probability

 finding n of binomial distribution given the probability from miscellaneous 12th cbse ncert mathematics

How many times must a man toss a fair coin so that the probability of having at least one head is more than 90%?

Let n be the required number of times the man must toss a fair coin so that the probability of having at least one head is more than 90%

Let X be the number of heads obtained when the coin is tossed n times.
Assuming X follows binomial distribution with
n=n
p = (½)
q = 1 – p = (½)

P[X=r] = nCr prq(n-r) , r = 0,1,2,...,n



disclaimer
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work

link to index of other miscellaneous problems on probability of cbse ncert 12th mathematics

index of more problems on baye's theorem for ncert cbse mathematics 

positive determinant from miscellaneous probability questions ncert cbse 12th mathematics

positive determinant and electronic assembly with two subsystems question from miscellaneous probability questions ncert cbse 12th mathematics

If each element of a second order determinant is either zero or one, what is the probability that the value of the determinant is positive?

There are 4 entries in the second order determinant ,each of which can be filled in two ways with either 0 or 1
Such determinants can be constructed in 24 ways ( (2^4) ways)
Therefore if S is the sample space n(S) = (2^4) = 16
Let E be the event that the value of the selected determinant is positive. 


Required probability = P(E) = [n(E)] / [n(S)] = 3 /16

=======================================

An electronic assembly consists of two subsystems, say, A and B. From previous testing procedures, the following probabilities are assumed to be known
P(A fails) = 0.2 P(B fails alone) = 0.15 P(A and B fail) = 0.15
Evaluate the following probabilities P(A fails|B has failed) ; P(A fails alone)

P(B fails) =P(A and B fail together ) + P(B fails alone) = 0.15+0.15 = 0.3

P(A fails|B has failed) = [P(A and B fail)] / [P(B fails ) ] = [0.15 / 0.30 =(1/2) = 0.5

P(A fails alone) = P(A fails) - P(A and B fail together ) = 0.2 – 0.15 = 0.05

disclaimer
There is no guarantee about the data/information on this site. You use the data/information at your own risk. You use the advertisements displayed on this page at your own risk.We are not responsible for the content of external internet sites. Some of the links may not work
 

Conic Sections in Precalculus: Equations, Graphs and Examples

  Conic Sections in Precalculus: Equations, Graphs, Examples and Practice Problems Conic sections are among the most useful topics in precal...